OpenStudy (anonymous):

A body of mass 150 kilograms is suspended from the ceiling with two wires that make an angle of 40.0 degrees witht he ceiling. What it the tension in each wire

OpenStudy (loinducoeur):

Draw a free body diagram first; it'll help a lot! And then, you can work out the tension by using: \[T _{2}\cos \theta _{2} = weight - T _{1}\cos \theta _{1}\]Does that make any sense to you? If it doesn't, I can show you some workings out and hopefully you could get the answer from that.